OpenStudy (anonymous):

Write the quotient in simplest form 2x+1/4x^2÷4x+2/x^2

OpenStudy (anonymous):

(2x+1/4x^2).(x^2/4x+2)=(2x+1)/4(4x+2)=(2x+1)/8(2x+1)=1/8

OpenStudy (anonymous):

[(2x+1)/4x^2]*[x^2/2(2x+1)]=1/8